A andB together can do a piece of work in 15days . IfA's one day work is 3/2 times one days work of B. Find in how many days can each do the work.?


Answer by stanbon(75887)

A and B together can do a piece of work in 15days . If A's one day work is 3/2 times one days work of B. Find in how many days can each do the work.?
-----
Let B's time to do the job alone be "x",
then A does the job in x/(3/2) = 2x/3 days
-----------------------------------------
Together rate:: 1/15 job/hr
A's rate:: 1/(2x/3) = 3/(2x) job/hr
B's rate:: 1/x job/hr
----
Equation::
rate + rate = together rate
1/x + 3/(2x) = 1/15
30 + 45 = 2x
x = 75/2 = 37.5 days (time for B to do the job alone)
2x/3 = (2/3)(37.5) = 25 days (time for A to do the job alone)
-------------
Cheers,
Stan H.
-------------